OpenStudy (anonymous):

-28.77 , -8.08, 8.0008, -0.5, 4.6 , -5.55, 10.00001, 10.01, 10.1 Convert to fraction as many as you can please and THANK YOU!! :DD

OpenStudy (anonymous):

-28.77 = -(2877/100) -8.08 = -(202/25) 8.0008 = (10001/1250) -0.5 = -(1/2) 4.6 = (23/5) -5.55 = -(111/20) 10.00001 = (1,000,001/100,000) 10.1 = (101/10) very easy to do if u have a graphing calculator

OpenStudy (anonymous):

Thank you . You save my sorry self thank you so much i owe you one

OpenStudy (anonymous):

out of all decimals which one is the largest